OpenStudy (anonymous):

1.2y-0.4y^2-0.3y-1.5y^2

OpenStudy (anonymous):

.9y-1.9y^2=y(.9-1.9y) If the equation is set to equal zero the solutions are 0,9/19

OpenStudy (anonymous):

when you set .9-1.9y=0 you get -1.9y=-.9 then y= -.9/-1.9=.9/1.9=9/19

OpenStudy (anonymous):

thank you